To avoid exciting at the PRM violin mode frequency, I have changed all of the filters relevant to the sensing matrix measurement from 628Hz to 580.1Hz. This includes notches in the LSC control loops, as well as the band pass filters in the lockins. I have not yet loaded the new filters, since arm locking is in progress.
I have loaded these new filters in. Manasa is still using the IFO for green stuff, so I can try out the PRMI measurement in a day or so. (Right now I have to make sure I understand my data, anyway.)